OpenStudy (anonymous):

Please Help. I have a homework online and I badly need the answer ASAP. The interest on an investment varies directly as the rate of interest. If the interest is $47 when the interest rate is 5% find the interest when the rate is 9.1%

OpenStudy (caozeyuan):

ok so this is a math problem, no finance

OpenStudy (caozeyuan):

all you have to do is find the relaionship between interes and the rate

OpenStudy (caozeyuan):

from the wording of the problem, it seems like interest is a funcion of rate so i= f(r)

OpenStudy (caozeyuan):

and the problem also says vary DIRECTLY, whihc tellls that f(r)=kr where k is some arbitrary constant

OpenStudy (caozeyuan):

since i=47 when r = 5% the constant k should be 47/5%= 940

OpenStudy (caozeyuan):

now r changes to 9.1% so i=940*9.1%=85.54

OpenStudy (caozeyuan):

so 85.54 is your answer

OpenStudy (anonymous):

Thank you! :)